An open flask sitting in a lab refrigerator looks empty, but it is actually filled with a mixture of gases called air. If the flask volume is 1.50 L, and the air is at standard temperature and pressure, how many gaseous molecules does the flask contain? number of gaseous molecules:
Answer #1

Molecules are calculated:

molecules = P * V * NA / R * T = 1 atm * 1.5 L * 6.02x10 ^ 23 molecules / mol / 0.082 atm * L / mol * K * 273 K = 4.03x10 ^ 22 air molecules.

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.